OpenStudy (anonymous):

Find the vertices and foci of the hyperbola with equation ((x+4)^2/9)-((y-5)^2/16)=1

OpenStudy (anonymous):

\[\frac{ (x+4)^2 }{ 9 }-\frac{ (y-5)^2 }{ 16 }=1\]

OpenStudy (anonymous):

lmfao

OpenStudy (antonio_xx2):

a^2 = 16, then a = ? answere=(8, -4) and (-2, -4)

OpenStudy (anonymous):

a=4

OpenStudy (anonymous):

Choices Vertices: (5, 0), (5, -8); Foci: (5, -8), (5, 0) Vertices: (5, -1), (5, -7); Foci: (5, -9), (5, 1) Vertices: (-1, 5), (-7, 5); Foci: (-9, 5), (1, 5) Vertices: (0, 5), (-8, 5); Foci: (-8, 5), (0, 5)

OpenStudy (anonymous):

b=3

OpenStudy (antonio_xx2):

Vertices: (5, 0), (5, -8); Foci: (5, -8), (5, 0)